# Globally defined Carroll symmetry of gravitational waves

### v2

@@ Abstract
The locally defined Carroll symmetry of a gravitational wave is extended to a globally defined one by switching to Brinkmann coordinates. Translations and Carroll boosts associated with two independent globally defined solutions of a Sturm-Liouville equation allow us to describe the motions. The Displacement Memory Effect arises for particular choices of the parameters which yield trajectories with zero momentum. The relation to the Schwarzian derivative is highlighted. We illustrate our general statements by the Pöschl-Teller profile.
@@ I Introduction
Braginsky and Thorne had suggested that gravitational waves might be observed by detecting the displacement of particles initially in rest, called later the Memory Effect BraTho . While experimental verification is still in the making, an insight can be gained by taking advantage of the Carroll symmetry Leblond ; SenGupta . Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen provide a simple description Sou73 ; Carroll4GW , however suffer of being regular only in finite intervals, requiring to glue them together. In this paper we show that such complications can be avoided by switching to Brinkmann (B) coordinates Bri by solving a matrix Sturm-Liouville equation, (II.3) below. The price to pay is, though, to get more complicated-looking expressions.
We illustrate our general theory by the Pöschl-Teller profile PTeller ; Plyushchay09 ; ChaKar , which is a very good analytical approximation of the widely studied Gaussian one. We follow mostly our review EZHRev which provides the reader also with further references.
Our notations are: $({\bm{X}},U,V)$ are Brinkmann coordinates and $({\bm{x}},u,v)$ are BJR coordinates on relativistic spacetime. $U$ is an affine parameter for geodesics. $\Theta^{Brink}={\bm{\xi}}{\partial}_{{\bm{X}}}+\eta{\partial}_{V}$ denote vector fields which generate the infinitesimal isometries on spacetime.
@@ II From Brinkmann to BJR and back
Plane gravitational waves are conveniently described by using globally defined Brinkmann coordinates Bri ; exactsol in terms of which the metric is,
@@
where ${\bm{X}}=(X^{1},\dots,X^{D})$ parametrizes the transverse plane, endowed with the flat Euclidean metric $d{\bm{X}}^{2}=\delta_{ij}\,dX^{i}dX^{j}$. $U$ and $V$ are light-cone coordinates and $(K_{ij})$ we call the profile is a symmetric $D\times D$ matrix whose entries depend only on $U$. Henceforth we focus our attention at $D=1$ or $D=2$ and assume, for simplicity, that is diagonal. In $D=2$ we take
@@
which yields a vacuum solution of the Einstein equations Bri ; exactsol . We consider “sandwich waves” BoPi89 , whose profile ${\mathcal{A}}(U)$ vanishes outside an interval $U_{i}<U<U_{f}$. The $V$-equation follows from the transverse ones EZHRev and will henceforth not be studied.
After the wave has passed, particles initially at rest exhibit the Velocity Memory Effect (VM) : they move with constant velocity. Under certain “quantization” conditions the outgoing velocity can vanish, though, and we get the Displacement Memory Effect (DM) ZelPol ; DM-1 ; DM-2 ; Sila-PLB , which might well play a rôle in future observations BraTho ; DM-2 .
Plane gravitational waves have long been known to have a $2D+1$ parameter symmetry group composed of $D+1$ translations, completed by $D$ (rather mysterious) transformations BoPiRo ; Sou73 . More recently Carroll4GW , this group was identified with the Carroll group Leblond ; SenGupta , as young Lévy-Leblond called it jokingly. Its algebraic structure is readily determined by switching to Baldwin et al (BJR) coordinates BaJe ; Rosen ; Sou73 which are however only local. This note sheds further light at the Carroll – Memory relation using globally defined Brinkmann coordinates.
@@
${P}^{\prime\prime}={\mathcal{A}}(U)P,\qquad(P^{T}){P}^{\prime}=(P^{T})^{\prime}P\,$ (II.3)
@@
for a $D{\times}D$ matrix $P(U)=\big(P_{ij}(U)\big)$ EZHRev . The BJR coordinates $({\bm{x}},u,v)$, are then obtained by
${{\bm{X}}}=P(u){\bm{x}},\qquad U=u,\qquad V=v-\frac{1}{4}{\bm{x}}\cdot\big(P^{T}{}P\big)^{\prime}{\bm{x}}\,$, (II.4)
in terms of which the metric is,
$g_{\mu\nu}dx^{\mu}dx^{\nu}=\big(P(u)^{T}P(u)\big)_{ij}(u)\,dx^{i}dx^{j}+2du\,dv\,.$ (II.5)
@@
Carroll symmetry consists, in BJR coordinates, of “vertical” ($v$) and transverse ($\bm{x}$}) translations with parameter $h$ and $\bf c$}, completed by Carroll boosts with parameter $\bf b$} Sou73 ; Carroll4GW ,
${\bm{x}}\to{\bm{x}}+{\bf c}+S(u)\,{\bf b}\,,\qquad u\to u,\qquad v\to v-{\bf b}\cdot{\bm{x}}-{{\scriptstyle{\frac{1}{2}}}}{\bf b}\cdot{}S(u)\,{\bf b}+h\,$, (II.6)
where the symmetric $2\times 2$ matrix
$S(u)=\int^{u}_{u_{0}}\!\left(P^{T}P\right)^{-1}\!(t)dt$ (II.7)
is referred to as the Souriau matrix Sou73 ; Carroll4GW ; EZHRev . Infinitesimal boosts parametrized with $b_{j}$ are,
$\theta^{BJR}=h\partial_{v}+c^{j}\partial_{j}+b_{j}(S^{ji}\partial_{i}-x^{j}\partial_{v})\,.$ (II.8)
@@
Finding the trajectories is then straightforward. Noether’s theorem provides us with $2D+1$ associated conserved quantities Sou73 ; Carroll4GW ,
${\bm{p}}={\bm{p}}_{0}=(P^{T}P)(u)\,{{\bm{x}}}^{\prime}(u),\qquad{\bf k}={\bf k}_{0}={\bm{x}}(u)-S(u)\,{\bm{p}}\,$, (II.9)
interpreted as conserved linear and boost-momentum, supplemented by $m={u}^{\prime}=1$. Then the conserved quantities determine the geodesic,
${\bm{x}}(u)={\bf k}_{0}+S(u)\,{\bm{p}}_{0}\,.$ (II.10)
@@
The subtlety comes from that the BJR coordinates are defined only in intervals $I_{k}=[u_{k-1},u_{k}]$ between the (mandatory Sou73 ; Carroll4GW ) zeros of the determinant,
$\det P(u_{k})=0\,$, (II.11)
@@
which play a fundamentally important rôle in our investigation, as we shall see. The BJR coordinates are thus only local : both the metric (II.5) and the Souriau matrix $S$ in (II.7) are singular at the junction points $u_{k}$, as will be illustrated in FIG.1 for the Pöschl - Teller profile. The singularity problem will be resolved in sect. III by switching to Brinkmann coordinates.
@@
Now we set $P^{T}=P$ and we consider $D=1$ (when $P$ is a scalar function) for simplicity, although our results and conclusions could be extended to non-diagonal profiles and to any dimension. Let us assume that $u<u_{1}$, the first zero of $\det(P)$. Requiring that our particles be in rest before the wave arrives,
${{\bm{x}}}(-\infty)=x_{0}{\quad\text{\small and}\quad}{{\bm{x}}}^{\prime}(-\infty)=0\,$, (II.12)
imply that the momentum vanishes in $I_{1}$,
${\bm{p}}=0\,.$ (II.13)
@@
The Souriau term is thus switched off from (II.10) and the transverse BJR “trajectory” is merely a fixed point.
${\bm{x}}(u)={\bm{x}}_{0}=\mathop{\rm const}\nolimits\,$, (II.14)
confirming, in BJR coordinates, the “no motion for Carroll” maxim Leblond . In contrast, the Brinkmann trajectory is non-trivial : (II.4) yields,
${\bm{X}}(U)=P(u)\,{\bm{x}}_{0}\,$, (II.15)
with $U=u$. The initial conditions (II.12) then require :
$P(-\infty)=1{\quad\text{\small and}\quad}P^{\prime}(-\infty)=0\,.$ (II.16)
@@
The Brinkmann trajectory, given by the $P$-matrix, may look quite complicated, depending on the Sturm-Liouville solution $P$ of (II.4) EZHRev . For appropriate “quantized” values of the wave parameter the wave zone contains an integer number of half-waves; then $P(u)$ tends to a constant matrix also for $u\to\infty$ : DM is obtained DM-1 ; DM-2 ; Sila-PLB .
@@
We emphasise that the investigations above are a priori valid only before the first zero of $\det{P}$ and should then be restarted until the next zero, and so on.
@@
Another remarkable consequence of (II.15) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from an $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin, and we get a caustic point BoPi89 ; EZHRev . The question will be further discussed in Sec.V and illustrated in FIG. 2.
@@ III Carroll symmetry in Brinkmann coordinates
Simple as they are, BJR coordinates (II.4) suffer from being valid only in intervals between subsequent zeros of $P$, and the adjacent expressions should be glued together – which may be laborious. A global approach can conveniently be given instead by using Brinkmann coordinates. For simplicity we restrict our attention at $D=1$ transverse dimension and start with a fixed interval $I_{k}$. A straightforward computation then shows that pulling back the BJR expression (II.6) by using (II.4) [or its infinitesimal form (II.8)] yields the Carroll generators written in Brinkmann coordinates,
@@
The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the the symmetry generators $\xi$ and $\eta$, mentioned in the Introduction. In Sec.V they will be spelt out explicitly for Pöschl - Teller .
The $P$-matrix and its derivative $P^{\prime}$ generate translations, whereas $PS$ and $(PS)^{\prime}$ generate Carroll boosts GenRelGrav . Their only nonzero commutator yields a vertical translation Blau24 ,
@@
$P$ thus plays a double role : by (II.15) it determines the B trajectory, whereas (III.1) says that it also generates translations. The B trajectory is obtained indeed from $x_{0}=\mathop{\rm const}\nolimits$ by a (generally $U$-dependent) translation.
The Brinkmann form (III.1) has an important advantage w.r.t. the BJR expression in (II.6) : multiplication by $P$ supresses the singularity of Souriau matrix $S$ Sou73 ; EZHRev  :
@@
is regular for all $U$. Thus (III.1) extends the Carroll symmetry (II.6) naturally from a fixed interval $I_{k}$ to the entire $U$-axis.
The remarkable fact is that both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.3) EZHRev ; Jibril ; Blau24 , as it can be verified by a direct calculation. The Wronskian is $W(P,Q)=P^{\prime}Q-Q^{\prime}P=-1$, therefore $P$ and $Q$ are independent. The expression (III.1) could also be verified independently, by checking the symmetry equation $L_{Y}g=0$ Torre ; GenRelGrav .
Contracting the Brinkmann metric (II.1) with (III.1) provides us with,
${\bm{p}}_{0}=P{\bm{X}}^{\prime}-\big(P\big)^{\prime}{\bm{X}}{\quad\text{\small and}\quad}{\bf k}_{0}=-Q{\bm{X}}^{\prime}+\big(Q\big)^{\prime}{\bm{X}}\,$, (III.4)
where ${\bm{p}}_{0}$ and ${\bf k}_{0}$ are, a priori, new constants of the motion. However expressing them in BJR terms, (II.4), the previous expressions (II.9) are recovered; as anticipated by our notation.
The Sturm-Liouville solutions $P$ and $Q$ play also yet another rôle, though : either combining (II.10) with (II.4) or solving (III.4) directly yields the general Brinkmann trajectory,
${{\bm{X}}(U)=P(U)\,{\bf k}_{0}+Q\,{\bm{p}}_{0}\,.}\;\;$ (III.5)
which is our second main result. When ${\bm{p}}_{0}=0$ cf. (II.13), then the $Q$ - term is switched off and we recover (II.15) with ${\bm{x}}_{0}={\bf k}_{0}$. For parameters which correspond to having an integer number of half-waves in the wavezone, we have also $P(+\infty)=\mathop{\rm const}\nolimits$ and we get DM.

### v3

@@ Abstract
The local Carroll symmetry of a gravitational wave found in Baldwin-Jeffery-Rosen coordinates is extended to a globally defined one by switching to Brinkmann coordinates. Two independent globally defined solutions of a Sturm-Liouville equation allow us to describe both the symmetries (translations and Carroll boosts) and the geodesic motions. One of them satisfies particular initial conditions which imply zero initial momentum, while the other does not. Pure displacement arises when the latter is turned off by requiring the momentum to vanish and when the wave parameters take, in addition, some particular values which correspond to having an integer half-wave number. The relation to the Schwarzian derivative is highlighted. We illustrate our general statements by the Pöschl-Teller profile.
@@ I Introduction
Braginsky and Thorne BraTho argued that gravitational waves might be observed by detecting the displacement of particles initially in rest, called later the Memory Effect. While experimental verification is still in the making, conceptual insight can be gained by taking advantage of the Carroll symmetry LeblondCar ; SenGupta . Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen provide a simple description Sou73 ; Carroll4GW , however are regular only in finite intervals, requiring to glue together the local results.
In this paper we show that such complications can be avoided by switching to globally defined Brinkmann (B) coordinates Bri ; DGH91 . The transcription requires to solve a matrix Sturm-Liouville equation, (II.6) below. The price to pay for globality is to get more complicated-looking expressions, though. The advantages and drawbacks of the two coordinate systems is discussed in the Conclusion.
We illustrate our general theory by the Pöschl-Teller profile PTeller ; Plyushchay09 ; ChaKar , which is a good analytical approximation of the widely studied Gaussian one. The book exactsol and our review EZHRev provide the reader with further references.
Our notations are: $({\bm{X}},U,V)$ are Brinkmann coordinates and $({\bm{x}},u,v)$ are BJR coordinates. $U=u$ is an affine parameter for geodesics. $\Theta^{Brink}={\bm{\xi}}{\partial}_{{\bm{X}}}+\eta{\partial}_{V}$ denotes collectively the vector fields which generate the infinitesimal isometries of spacetime. Its explicit form will be spelt out in (III.1).
@@ II From Brinkmann to BJR and back
Plane gravitational waves are conveniently described by using globally defined Brinkmann coordinates Bri ; DGH91 ; exactsol in terms of which the metric is,
@@
where the ${\bm{X}}=(X^{1},\dots,X^{D})$ are coordinates of the transverse plane with the flat Euclidean metric $d{\bm{X}}^{2}=\delta_{ij}\,dX^{i}dX^{j}$. $U$ and $V$ are light-cone coordinates related to the usual relativistic coordinates according to $U=\frac{1}{\sqrt{2}}(Z-T)$ and $V=\frac{1}{\sqrt{2}}(Z+T)$. The profile $(K_{ij})$ is a symmetric $D\times D$ matrix whose entries depend only on $U$. Henceforth we focus our attention at $D=1$ or $D=2$ and assume, for simplicity, that the wave is linearly polarized so that $(K_{ij})$ is diagonal. Thus in $D=2$, which is our main interest in this paper,
@@
which yields a vacuum solution of the Einstein equations for any ${\mathcal{A}}(U)$ Bri ; exactsol ; DGH91 ; EZHRev . We consider only “sandwich waves” BoPi89 , whose profile ${\mathcal{A}}(U)$ vanishes outside an interval $U_{i}<U<U_{f}$. Choosing $U$ as an affine parameter, the equations of motion of the transverse coordinates are of the Sturm-Liouville form,
$\dfrac{d^{2}{\bm{X}}}{dU^{2}}-\left(\begin{array}[]{lr}{\mathcal{A}}&0\\ 0&-{\mathcal{A}}\end{array}\right)\,{\bm{X}}=0\,.$ (II.3)
The geodesic equations have an additional constant of the motion called the Jacobi invariant exactsol ,
$e=\frac{1}{2}\,g_{\mu\nu}\,\frac{d{X}^{\mu}}{dU}\frac{d{X}^{\nu}}{dU}\,.$ (II.4)
The transverse equations (II.3) do not depend on $e$. In this paper we limit our attention at the lightlike case $e=0$ for which the “vertical” coordinate $V$ satisfies
$\dfrac{d^{2}V}{dU^{2}}+\dfrac{1}{2}\displaystyle{\frac{d{\mathcal{A}}}{dU\,}}\Big((X^{1})^{2}-(X^{2})^{2}\Big)+2{\mathcal{A}}\Big(X^{1}\dfrac{dX^{1}}{dU\,}-X^{2}\dfrac{dX^{2}}{dU\,}\Big)=0\,.$ (II.5)
It corresponds to the horizontal lift of (II.3) and can be solved once the solutions of the latter are known. Henceforth we focus our attention at (II.3). The timelike (massive) case, $e<0$, is discussed, e.g. in EDAHK ; DM-1 . The equations above constitute a model for particle dynamics in a gravitational wave including flyby ZelPol .
After the wave has passed, particles initially at rest exhibit the Velocity Memory Effect (VM) Ehlers ; AiBalasin ; EZHRev  : outside the Wavezone, they move with constant velocity. Under certain “quantization” conditions the outgoing velocity can vanish, though, and we get the Displacement Memory Effect (DM) ZelPol ; DM-1 ; DM-2 ; Sila-PLB .
Plane gravitational waves have long been known to have a $2D+1$ parameter symmetry group composed of $D+1$ translations, completed by $D$ (rather mysterious) transformations BoPiRo ; Sou73 . More recently Carroll4GW , this group was identified with the Carroll group LeblondCar ; SenGupta , as young Lévy-Leblond called it jokingly. Its algebraic structure is readily determined by switching to Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen ; Sou73 which are however only local. This note sheds further light at the Carroll — Memory relation using globally defined Brinkmann coordinates.
@@
${P}^{\prime\prime}={\mathcal{A}}(U)P,\qquad(P^{T}){\,}{P}^{\prime}=(P^{T})^{\prime}{\,}P\,$ (II.6)
@@
for a $D{\times}D$ matrix $P(U)=\big(P_{ij}(U)\big)$ EZHRev . The prime here denotes $d/du\equiv d/dU$. BJR coordinates $({\bm{x}},u,v)$ are then obtained by
${{\bm{X}}}=P(u){\bm{x}},\qquad U=u,\qquad V=v-\frac{1}{4}{\bm{x}}\cdot\big(P^{T}{}P\big)^{\prime}{\bm{x}}\,$, (II.7)
in terms of which the metric takes the BJR form,
$g_{\mu\nu}dx^{\mu}dx^{\nu}=\left(P(u)^{T}P(u)\right)_{ij}(u)\,dx^{i}dx^{j}+2du\,dv\,.$ (II.8)
@@
Carroll symmetry written in BJR coordinates is composed of “vertical” ($v$) and transverse ($\bm{x}$}) translations with parameters $h$ and $\bf c$}, respectively, completed by Carroll boosts with parameter $\bf b$} Sou73 ; Carroll4GW ,
${\bm{x}}\to{\bm{x}}+{\bf c}+S(u)\,{\bf b}\,,\qquad u\to u,\qquad v\to v-{\bf b}\cdot{\bm{x}}-{{\scriptstyle{\frac{1}{2}}}}{\bf b}\cdot{}S(u)\,{\bf b}+h\,.$ (II.9)
The symmetric $2\times 2$ matrix
$S(u)=\int^{u}_{u_{0}}\!\left(P^{T}P\right)^{-1}\!(t)dt$ (II.10)
here is referred to as the Souriau matrix Sou73 ; Carroll4GW ; EZHRev . $S$ is related to but is slightly different from the time redefinition considered, in the isotropic case, by Arnold Arnold ; ZurabNote ; Aniso ; ZZH . Details will be as discussed elsewhere ZEHPR .
The infinitesimal symmetries are, with some abuse of notation,
$\theta^{BJR}=h\partial_{v}+c^{j}\partial_{j}+b_{j}\big(S^{ji}\partial_{i}-x^{j}\partial_{v}\big)\,.$ (II.11)
The covariantly constant Killing vector ${\partial}_{v}$ here generates “vertical” translations – a hallmark of Brinkmann metrics Bri ; DGH91 . The $c_{i}$ generate translations and the $b_{j}$ generate Carroll boosts. For $u=u_{0}$, the integral (II.10) vanishes and we recover the usual Carroll action in LeblondCar .
@@
Finding the geodesic trajectories is then straightforward by using the symmetry. Noether’s theorem provides us with associated conserved quantities Sou73 ; Carroll4GW ,
${\bm{p}}={\bm{p}}_{0}=(P^{T}P)(u)\,{{\bm{x}}}^{\prime}(u),\qquad{\bf k}={\bf k}_{0}={\bm{x}}(u)-S(u)\,{\bm{p}}\,$, (II.12)
interpreted as conserved linear and boost-momentum, respectively, supplemented by the non-relativistic mass that we scale to $1$. It is generated by the Killing vector ${\partial}_{v}$. The conserved quantities determine the geodesic trajctories  [^4] A non-vanishing Jacobi invariant $e\neq 0$ would merely add a linear term, $eu$ to the r.h.s., EDAHK ; DM-1 .,
${\bm{x}}(u)={\bf k}_{0}+S(u)\,{\bm{p}}_{0}\,,\quad v(u)=-\frac{1}{2}{\bm{x}}.{\bm{p}}+v_{0}\,.$ (II.13)
@@
So far so good. However a subtlety comes from that the BJR coordinates are defined only in intervals $I_{k}=[u_{k-1},u_{k}]$ between mandatory Sou73 ; Carroll4GW zeros of the determinant [^5] In $D=1$, $P(U)$ is just a scalar function.,
$\det P(u_{k})=0\,.$ (II.14)
@@
The matrix $P$ plays a fundamentally important rôle in our investigation, as we shall see. Both the metric (II.8) and the Souriau matrix $S$ (II.10) are singular at the junction points $u_{k}$, as is manifest in FIG.2 for the Pöschl - Teller profile. The singularity will be removed by switching to Brinkmann coordinates.
Although our results could be extended to non-diagonal profiles and to any dimension, henceforth we focus our attention at the diagonal case, $P^{T}=P$, and to D=2D=2i.e.,  to the physically relevant $4$-dimensional spacetime. Occasionally, we consider also toy examples in $D=1$, though.
@@
Let us assume that $u<u_{1}$, the first zero of $\det(P)$. Requiring that our particles be in rest before the wave arrives,
${{\bm{x}}}(-\infty)=x_{0}{\quad\text{\small and}\quad}{{\bm{x}}}^{\prime}(-\infty)=0\,$, (II.15)
implies that the momentum vanishes in $I_{1}$,
${\bm{p}}=0\,.$ (II.16)
@@
The Souriau term is thus switched off from (II.13) and the transverse BJR “trajectory” is, for vanishing Jacobi invariant, $e=0$ in (II.4), merely a fixed point  [^6] A non-vanishing Jacobi in $e\neq 0$ would merely shift $v(u)$ by a linear-in-$u$ term EDAHK ; DM-1 .,
${\bm{x}}(u)={\bm{x}}_{0}=\mathop{\rm const}\nolimits\,,\qquad v(u)=v_{0}=\mathop{\rm const}\nolimits\,$, (II.17)
consistently with (II.13). Eqn (II.17) confirms the “no motion for Carroll” wisdom LeblondCar   [^7] Motion is possible for multi-particle systems MultiCarroll , and for planar models with “exotic” double central extension when the particles follow the Hall law MARSOT ; ExoPeierls .. However, unlike as for BJR, the (lightlike) Brinkmann trajectory is non-trivial : (II.7) yields,
${\bm{X}}(U)=P(u)\,{\bm{x}}_{0}\,,\qquad V(U)=-\frac{1}{2}{\bm{X}}.{\bm{X}}^{\prime}+V_{0}\,$, (II.18)
with $U=u$ and $V_{0}=v_{0}$. The initial conditions (II.15) then require :
$P(-\infty)={\rm Id}{\quad\text{\small and}\quad}P^{\prime}(-\infty)=0\,.$ (II.19)
@@
The Brinkmann trajectory, given by the action of the $P$-matrix, (II.18), may look indeed quite complicated, depending on the Sturm-Liouville solution $P$ of (II.6) EZHRev . For randomly chosen wave parameters we get VM Ehlers ; AiBalasin ; EZHRev  : the particles fly apart with constant velocity. For appropriately “quantized” values of the wave parameter, though, the wave zone contains an integer number of half-waves; $P(u)$ then tends to a constant matrix also for $u=U\to\infty$ and DM is obtained DM-1 ; DM-2 ; Sila-PLB .
@@
We emphasise, though, that the investigations above are a priori valid only before the first zero of $\det{P}$ and should then be restarted until the next zero, and so on, as it will be seen in sec.V for Pöschl - Teller .
@@
Another remarkable consequence of (II.18) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin : we get a caustic point BoPi89 ; EZHRev , as it will be illustrated in FIG. 3.
@@ III Carroll symmetry in Brinkmann coordinates
Simple as they are, BJR coordinates (II.7) suffer from being valid only in intervals between subsequent zeros of $det(P)$. The adjacent expressions should be glued together – which may be laborious. A global approach can conveniently be given instead by using Brinkmann coordinates. For simplicity we restrict our attention at $D=1$ transverse dimension and start with a fixed interval $I_{k}$. A straightforward computation then shows that pulling back the BJR expression (II.9) by (II.7) [or its infinitesimal form (II.11)] yields the Carroll generators written in Brinkmann coordinates,
@@
which could also be verified independently, by checking the equation $L_{Y}g=0$ Torre ; GenRelGrav . The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the symmetry generators $\xi$ and $\eta$ mentioned in the Introduction. They will be spelt out explicitly for Pöschl - Teller in Sec.V. The $P$-matrix and its derivative $P^{\prime}$ generate translations, whereas $Q=PS$ and $Q^{\prime}=(PS)^{\prime}$ generate Carroll boosts GenRelGrav . ${\partial}_{V}$ generates the non-relativistic mass scaled to $1$, as does for BJR.
The only nonzero commutator yields a vertical translation Blau24 ,
@@
$P(U)$ thus plays a double role : for DM parameters it determines the complicated-looking Brinkmann trajectory in (II.18), whereas by (III.1) it also generates translations. For non-DM parameters, the situation is even more complicated, as will be seen in (III.6).
The Brinkmann form (III.1) has an important advantage w.r.t. the BJR expression (II.9) : left multiplication by $P$ removes the singularity of the Souriau matrix $S$ Sou73 ; EZHRev  :
@@
is regular for all $U$. In conclusion, switching form BJR to Brinkmann, (III.1),extends the Carroll symmetry (II.9) from an interval $I_{k}$ to the entire $U$-axis. A remarkable observation says that both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.6),
${Q}^{\prime\prime}={\mathcal{A}}(U)Q,\qquad(Q^{T}){\,}{Q}^{\prime}=(Q^{T})^{\prime}{\,}Q\,$, (III.4)
as it can also be verified by a direct calculation. $P$ and $Q$ have different initial conditions, though EZHRev ; Jibril ; Blau24 . The two solutions are independent when their Wronskian, $W(P,Q)=P^{\prime}Q-Q^{\prime}P$, does not vanish.
Contracting the Brinkmann metric (II.1) with the Carroll generators (III.1) provides us with,
${\bm{p}}_{0}=P{\bm{X}}^{\prime}-\big(P\big)^{\prime}{\bm{X}}{\quad\text{\small and}\quad}{\bf k}_{0}=-Q{\bm{X}}^{\prime}+\big(Q\big)^{\prime}{\bm{X}}\,$, (III.5)
where ${\bm{p}}_{0}$ and ${\bf k}_{0}$ are, a priori, new constants of the motion. However expressing them in BJR terms using (II.7), the previous expressions (II.12) are recovered, as anticipated by our notation.
The Sturm-Liouville solutions $P$ and $Q$ play yet another rôle, though : either combining (II.13) with (II.7) or solving (III.5) directly yields the general Brinkmann trajectory,
${{\bm{X}}(U)=P(U)\,{\bf k}_{0}+Q(U)\,{\bm{p}}_{0}\,.}\;\;$ (III.6)
Remember that $P$ satisfies the initial conditions (II.19).
For vanishing momentum ${\bm{p}}_{0}=0$, cf. (II.16), the VM term $Q$ is switched off and we recover (II.18) with ${\bm{x}}_{0}={\bf k}_{0}$. The corresponding Brinkmann trajectory may or may not be DM yet, though, depending on the wave profile. For parameters for which we have an integer number of half-waves in the wavezone, we do get DM : $P(+\infty)=\mathop{\rm const}\nolimits$, though DM-1 ; DM-2 ; Sila-PLB . The case of nonzero initial momentum, ${\bm{p}}_{0}\neq 0$, will be studied in ZEHPR .